OpenStudy (anonymous):

Which of the following is not true? 2 and 5 are the only prime factors of twenty. The prime factorization of forty-six is 2 times twenty-three. 3 is the only prime factor of twenty-seven. The prime factorization of twenty-four is 2 squared times 3 squared

OpenStudy (anonymous):

\[20=2^{2}*5\]\[46=2*23\]\[27=3^{3}\]\[24=2^{3}*3\] Looks like the last one of yours is not true. THe prime factorization of twenty-four is two cubed times three.

jhonyy9 (jhonyy9):

the last is 4*9=36